" Year Stock X Stock Y 1 9.00% -6.00% 2 -8.00% 11.00% 3 18.00% 2.00% 4...

"

Year

Stock X

Stock Y

1

9.00%

-6.00%

2

-8.00%

11.00%

3

18.00%

2.00%

4

3.00%

5.00%

What is the covariance of the returns of Stock X with the returns of Stock Y?"

Round to 6 decimal places please.
Answer #1

Covariance =

Cov(x,y) = Σ ((xi – x) * (yi – y)) / N

here xj= return of stock x

x = mean of return of stockx in jth interval

yj = return of stock y

y = mean of return of stock y in jth interval

N = number of intervals , i.e. 4

Step1 : finding mean

mean = sum of stock returns/number of intervals

mean of x = (0.09-0.08+0.18+.03)/4 = 0.055

mean of y = (-0.06+0.11+0.02+0.05)/4 = 0.03

Step2 : Finding difference between each value and mean price

Year (A) Stock X (B) return of X
( C)

x - mean x

(D)

Stock Y ( E) return of Y (F)

y - mean y

(G)

(D)*(G)
1 9.00% 0.09 0.035 -6.00% -0.06 -0.09 -0.00315
2 -8.00% -0.08 -0.135 11.00% 0.11 0.08 -0.0108
3 18.00% 0.18 0.125 2.00% 0.02 -0.01 -0.00125
4 3.00% 0.03 -0.025 5.00% 0.05 0.02 -0.0005
mean x 0.055 mean y 0.03 sum = -0.0157


= -0.0157

Step3: Dividing by N

Covariance = -0.0157/4

= -0.003925 ( rounded to 6 decimal places)